# Decoding and Engineering the Phytobiome Communication for Smart Agriculture

### v1

@@ IV Challenges and Open Research Issues
As detailed in Section III, targeted delivery of agrochemicals in the phytobiome is a promising research direction, with ongoing MC research using magnetic nanoparticles offering a potential solution [14]. In addition, phytobiome-specific methods are needed to localize target cells with NMs. While the research on IoBNT is currently at an experimental stage, its adaptation to agricultural settings offers a lower-cost, more accessible alternative to biomedical IoBNT scenarios. On the other hand, plant stress monitoring via electrophysiological signals and ML/AI is at a more advanced stage in both academia and industry. However, it is imperative to investigate how the intricate structure of phytobiome communication in field conditions affects the accuracy of stress detection.
@@ V Industrial and Agricultural Outlook
For the last 30 years, increased agricultural efficiency has been driven partly by improved use of data in a collection of technologies often termed precision agriculture (PA)[15]. In PA, farmers collate information from the environment, satellites, drones, weather stations, soil moisture sensors, and cameras mounted on tractors. These data are then used to respond in a more precise way to local conditions. However, the future of agriculture lies in smart agriculture, which goes beyond PA by incorporating advanced technologies such as the IoBNT, ML/AI, and robotics to create a more interconnected, automated, and responsive farming system.
@@ VI Conclusion
In this paper, we explore the complex phytobiome communication and its use in smart agriculture. We propose a multi-scale communication framework to decode the interactions between plants and associated organisms in their environment through molecular and electrophysiological communication signals. This framework offers a holistic modeling approach across different scales, facilitating the understanding of the phytobiome as a communication network. We leverage this approach to engineer the phytobiome communication with novel smart agriculture applications. Notably, the fully autonomous integration of smart irrigation and targeted agrochemical delivery via IoBNT, coupled with the diagnosis of plant stress using ML/AI on electrophysiological signals, holds the potential to eliminate human errors in sustainable, eco-friendly, and efficient agriculture.
@@ Acknowledgements
The authors would like to thank Magdalena Janicka for the artwork in Fig. 1. The work of Fatih Gulec was supported by the MSCA YUFE Postdoctoral Fellowship under Horizon Europe’s MSCA COFUND program. The work of Andrew W. Eckford was supported by a Discovery grant from NSERC.
@@
Fatih Gulec [M] received his Ph.D. degree in Electronics and Communication Engineering from Izmir Institute of Technology, Turkey, in 2021. He is currently an MSCA YUFE Postdoctoral Fellow with the Department of Computer Science and Electronic Engineering at the University of Essex, UK. His research interests include molecular communications, computational biology, and plant electrophysiology/bioacoustics. He was awarded a short-term DAAD Postdoctoral Fellowship in 2021, the MSCA YUFE Fellowship in 2024, and the 2022 Doctoral Thesis Award from the IEEE Turkey Section.

### v2

@@ IV Challenges and Open Research Issues
As detailed in Section III, targeted delivery of agrochemicals in the phytobiome is a promising research direction, with ongoing MC research using magnetic nanoparticles offering a potential solution [wicke2021experimental]. In addition, phytobiome-specific methods are needed to localize target cells with NMs. While the research on IoBNT is currently at an experimental stage, its adaptation to agricultural settings offers a lower-cost, more accessible alternative to biomedical IoBNT scenarios. On the other hand, plant stress monitoring via electrophysiological signals and ML/AI is at a more advanced stage in both academia and industry. However, it is imperative to investigate how the intricate structure of phytobiome communication in field conditions affects the accuracy of stress detection.
@@ V Industrial and Agricultural Outlook
For the last 30 years, increased agricultural efficiency has been driven partly by improved use of data in a collection of technologies often termed precision agriculture (PA)[sishodia2020applications]. In PA, farmers collate information from the environment, satellites, drones, weather stations, soil moisture sensors, and cameras mounted on tractors. These data are then used to respond in a more precise way to local conditions. However, the future of agriculture lies in smart agriculture, which goes beyond PA by incorporating advanced technologies such as the IoBNT, ML/AI, and robotics to create a more interconnected, automated, and responsive farming system.
@@ VI Conclusion
In this paper, we explore the complex phytobiome communication and its use in smart agriculture. We propose a multi-scale communication framework to decode the interactions between plants and associated organisms in their environment through molecular and electrophysiological communication signals. This framework facilitates the understanding of the phytobiome as a communication network, and we review our experimental results to illustrate how it can model electrophysiological signals via MC. We then leverage this framework to engineer phytobiome communication with novel smart agriculture applications. Notably, the fully autonomous integration of smart irrigation and targeted agrochemical delivery via IoBNT, coupled with ML/AI-based diagnosis of plant stress from electrophysiological signals, holds the potential to eliminate human errors in sustainable, eco-friendly, and efficient agriculture. A practical near-term direction is to apply our framework to increase the accuracy of electrophysiology-based stress monitoring by investigating how phytobiome communication under field conditions affects signal patterns.
@@ Acknowledgements
The authors would like to thank Magdalena Janicka for the artwork in Fig. 1. Fatih Gulec was supported by EU Horizon Europe under the Marie Skłodowska-Curie COFUND grant No 101081327 YUFE4Postdocs. The work of Andrew W. Eckford was supported by a Discovery grant from NSERC.
@@
Fatih Gulec [M] received his Ph.D. degree in Electronics and Communication Engineering from Izmir Institute of Technology, Turkey, in 2021. He is currently an MSCA YUFE Postdoctoral Fellow with the Department of Computer Science and Electronic Engineering at the University of Essex, UK. His research interests include molecular communications, mathematical biology, and plant electrophysiology/bioacoustics. He was awarded a short-term DAAD Postdoctoral Fellowship in 2021, the MSCA YUFE Fellowship in 2024, and the 2022 Doctoral Thesis Award from the IEEE Turkey Section.